# DrugAgent: Explainable Drug Repurposing Agent with Large Language Model-based Reasoning

### v2

@@ Abstract
Drug repurposing, the identification of new therapeutic applications for existing drugs, offers a promising approach to accelerate drug development. This paper introduces a novel multi-agent framework that enhances the drug repurposing process by leveraging state-of-the-art machine learning techniques and integrating diverse knowledge sources. Our framework consists of three specialized agents: an AI Agent that develops robust drug-target interaction (DTI) models; a Knowledge Graph Agent that extracts DTIs from established databases such as DGIdb, DrugBank, CTD, and STITCH; and a Search Agent that engages with biomedical literature to annotate and verify computational predictions. By synthesizing outputs from these agents, our system harnesses a wide array of data sources to propose viable repurposing candidates. Preliminary results demonstrate the framework’s potential to not only predict drug-disease interactions but also to reduce the time and costs associated with traditional drug discovery methods. Our approach outperforms existing methods in predicting drug repurposing potential while providing interpretable results. This work highlights the scalability of multi-agent systems in biomedical research and their capacity to drive innovation in drug repurposing. The framework presented in this paper offers a promising pathway towards more efficient and cost-effective drug discovery processes, paving the way for accelerated therapeutic developments. Code is available https://anonymous.4open.science/r/DrugRepurposeLLM-E0B5/.
@@ 1 Introduction
Our framework includes three primary agents: the AI agent, the Knowledge Graph (KG) Agent, and the Search Agent. The AI agent employs the DeepPurpose package (Huang et al., 2020) to develop and refine drug-target interaction (DTI) models. The knowledge graph Agent utilizes the drug-gene interaction database (DGIdb) (Cannon et al., 2024), DrugBank (Knox et al., 2024), Comparative Toxicogenomics Database (CTD) (Davis et al., 2023), and Search Tool for Interactions of Chemicals (STITCH) (Kuhn et al., 2007) to extract and synthesize information on DTIs. The search Agent engages with biomedical literature, leveraging large language models for automated data labeling and validation, a process resource-intensive and costly.
@@ 2 Related Works
Machine Learning in Drug Repurposing. Machine learning techniques have been applied to drug repurposing, demonstrating significant potential in predicting drug-disease interactions (Issa et al., 2021). Similarly, our AI Agent leverages the DeepPurpose toolkit, which has been extensively validated for its efficiency in drug-target interaction (DTI) predictions (Huang et al., 2020).
@@ 3.1 Overview of DrugAgent
To process natural language inputs and generate responses that are coherent and contextually appropriate, each agent utilizes large language models. The system’s reasoning capabilities are enhanced by incorporating methodologies that allow for step-by-step problem-solving and decision-making.
Following the reasoning process, the system is capable of taking actions such as calculating repurposing scores, performing literature-based analyses, and querying knowledge graphs for novel drug-disease associations. By integrating this information through a weighted average approach, the system simulates a highly knowledgeable drug repurposing researcher.
@@ 3.2 Agent Roles and Responsibilities
The DrugAgent framework integrates a diverse array of specialized agents, each employing the ReAct (Yao et al., 2022) and LEAST-TO-MOST (Zhou et al., 2022) reasoning methods to meticulously plan their actions. Through the use of advanced search capabilities, access to specialist models, and indexing in databases, these agents can execute a wide range of tasks effectively. Below, we delve into the specific roles and responsibilities assigned to each agent within the system.
@@ 3.3 AI Agent
The AI Agent’s predictions are continually refined through an iterative training and validation process, using cross-validation techniques to ensure accuracy and robustness.
@@ 3.5 Search Agent: Information Extraction from Biomedical Literature
This simplified approach allows for rapid information gathering from publicly available sources. However, it is important to note that this method relies on the quality and relevance of Google search results, and does not directly access or analyze full scientific papers or curated databases. As such, it serves as a preliminary screening tool rather than a comprehensive literature review system.
@@ 3.6 Calling External Tools
GPT supports calling external tools (e.g., function, database retrieval) to leverage external knowledge and enhance its capability. Specifically, suppose we have a couple of toolkits, GPT’s API can automatically detect which tool to use, which serves as glue to connect large language models to external tools. Our system integrates a variety of external data sources and predictive AI models to support the agents’ functions.
@@
  • Comparative Toxicogenomics Database (CTD): The CTD (Davis et al., 2023) is a robust, manually curated database that provides information about chemical–gene/protein interactions, chemical–disease, and gene-disease relationships. It is valuable for understanding how environmental exposures affect human health, integrating data from various species and linking chemicals, genes, diseases, phenotypes, and pathways (Chang et al., 2019; Wu et al., 2022a).
@@
We utilize DeepPurpose (Huang et al., 2020) for the AI Agent. DeepPurpose is a comprehensive and extensible deep learning library for DTI prediction. It integrates multiple state-of-the-art models and datasets, allowing researchers to easily implement and compare various deep learning approaches for drug discovery and repurposing. DeepPurpose facilitates the application of AI in drug development by providing a unified framework for different drug and protein encoding methods.
@@ 3.7 Drug Repurposing Agent
The predictions from the AI agent are cross-referenced and enriched with data from the Knowledge Graph Agent and validated against findings from the Search Agent. This integrated workflow allows for a dynamic updating mechanism, wherein feedback from the literature and knowledge graph continually refines the predictions. The entire process is supported by a central data hub that manages data consistency and real-time updates, ensuring that all agents have access to the latest and most accurate data available.
@@
$S_{merged}=\alpha S_{ml}+\beta S_{search}+(1-\alpha-\beta)S_{kg},$
where $S_{merged}$ is the merged DTI score, $S_{ml}$ is the machine learning DTI score, $S_{search}$ is the search-based DTI score, $S_{kg}$ is the knowledge graph DTI score, $\alpha$ is the weight for the machine learning score, $\beta$ is the weight for the search score, and $(1-\alpha-\beta)$ is the weight for the knowledge graph score. Note that $10<\alpha<$ 1, $10<\beta<$ 1, and $\alpha+\beta<1$.
@@ 3.8 Workflow
Step 6: Delivery of Solution. The final output, which includes the merged DTI score along with the individual scores from each method, is delivered to the user. This comprehensive result provides not only the final prediction but also insights into how different methods contribute to the overall score, enhancing the user’s understanding of the DTI prediction. This structured workflow ensures that our multi-agent DTI prediction system effectively combines multiple analytical approaches, offering a robust and multi-faceted assessment of potential DTIs.
@@ 4.1 Quantitative Results: Performance Comparison of DrugAgent and GPT-4
Table 1: Performance comparison between DrugAgent and GPT-4 for different datasets. We report the average results of 5 independent runs and the corresponding standard deviations (in brackets). For each metric, we highlight the best method in bold. We marked the metrics where DrugAgent is significantly better than GPT-4 (pass the t-test, i.e., p-value$<$0.05) using “*”.
@@
DrugAgent outperformed GPT-4 across all examined metrics. Regarding prediction accuracy, DrugAgent achieved a Mean Squared Error (MSE) of 1.836, significantly lower than GPT-4’s 13.420, indicating superior overall predictive power. The Mean Absolute Percentage Error (MAPE) and Mean Absolute Error (MAE) metrics further confirm this superiority, with DrugAgent achieving values of 0.134 and 1.081, respectively, compared to GPT-4’s 0.320 and 3.350.
@@
These results collectively showcase DrugAgent’s significantly improved predictive capabilities for BindingDB PKD scores compared to GPT-4, which is crucial for accurate binding affinity predictions in drug discovery and molecular interaction studies.
@@
DrugAgent achieves its superior performance with a runtime of 5.000s, which is efficient considering the complexity of its multi-agent architecture. In comparison, GPT-4’s runtime of 0.297s is faster, but at the cost of significantly reduced accuracy.
@@
This comprehensive analysis highlights the synergistic interaction between the AI, KG, and Search Agents in our model architecture for accurate PKD score prediction. The results validate our design choices and provide valuable insights for future improvements and optimizations of DrugAgent. They particularly emphasize the potential for enhancing the integration and capabilities of each component to achieve even more accurate PKD score predictions in the context of the BindingDB database, while considering the balance between performance, computational resources, and cost.
@@ 4.2 Case Study
This study presents three case studies analyzing the drug repurposing potential of Topotecan for different proteins: TOP1, SLFN11, and SLC26A4. These cases represent a spectrum from known strong interactions to potentially novel connections, allowing for a comprehensive evaluation of our multi-agent system’s capabilities. In each case, the Coordinator Agent decomposed the complex problem into three subproblems, each handled by a specialized agent. The AI Agent predicted interaction potential based on structural properties, the Search Agent analyzed evidence from scientific literature, and the Knowledge Graph Agent evaluated connections in existing databases.
@@
Case 1 examined the interaction between Topotecan and TOP1, a known strong interaction. The system calculated a final score of 11.51, confirming the established relationship. The high AI Agent score (7.65) and Knowledge Graph Agent score (1.0) aligned with the known mechanism of Topotecan as a TOP1 inhibitor, while the relatively low Search Agent score (0.27) likely reflected the well-established nature of this interaction not requiring extensive new studies.
@@
Case 2 investigated the less understood but potentially relevant interaction between Topotecan and SLFN11. The system yielded a final score of 10.30, suggesting a noteworthy relationship. The high AI Agent score (7.36) indicated structural compatibility, while the moderate Search Agent (0.33) and Knowledge Graph Agent (0.72) scores reflected some existing evidence and established connections, albeit less comprehensive than in Case 1.
@@
Case 3 explored an unlikely interaction between Topotecan and SLC26A4, resulting in a final score of 9.92. Despite a high AI Agent score (7.61) suggesting structural compatibility, the very low Search Agent score (0.00) indicated minimal literature evidence. The moderate Knowledge Graph Agent score (0.72) suggested some indirect connections, highlighting the system’s ability to detect potential novel interactions while also recognizing the need for careful interpretation.
@@
These case studies demonstrated the multi-agent system’s ability to handle diverse scenarios, from known interactions to potential discoveries. The system effectively integrated various data sources and analytical methods, providing interpretable results with detailed reasoning processes. Importantly, it showed the capacity to recognize its own limitations, particularly when predictions conflicted with existing evidence.
@@
The system offered practical insights for drug repurposing research across different levels of prior knowledge. In Case 1, it confirmed a well-established interaction. In Case 2, it suggested a potentially relevant interaction that warrants further investigation. In Case 3, it identified a possible novel interaction while appropriately flagging the low literature evidence, thus highlighting an area requiring careful experimental validation.
@@
This comprehensive approach not only validates known interactions but also suggests potential new avenues for research. By providing nuanced analyses that consider multiple factors, the system demonstrates its value in accelerating and refining the drug repurposing process. Future work could focus on experimental validation of the system’s predictions, particularly for cases like Topotecan-SLFN11 where moderate evidence suggests a promising avenue for investigation.
@@ 5 Discussion
Our study presents a novel multi-agent system for drug repurposing that integrates machine learning, knowledge graphs, and literature search. This approach offers more robust predictions by leveraging diverse data sources and analytical methods. The system’s strength lies in its collaborative approach, which combines each agent’s specialized capabilities to evaluate complex DTIs comprehensively. The weighted integration method allows for flexible adjustment of different prediction methods, enhancing overall accuracy. However, limitations exist. The system still relies on human expertise for initial setup, limiting its scalability. It also lacks autonomous knowledge updating capabilities to keep pace with rapidly evolving pharmacological research. Furthermore, the current model does not adequately account for individual patient characteristics or drug combination effects.
@@
Automating the preprocessing of complex data types, such as scRNA-seq data, is another important area for improvement. This involves implementing multiple preprocessing functions including Imputation (Inoue et al., 2024a; Inoue, 2024), Quality Control (McCarthy et al., 2017; Lu et al., 2023), and batch effect correction (Li et al., 2020; Fu et al., 2024; Haghverdi et al., 2018). Developing an optimization framework to automatically select and apply the most appropriate preprocessing methods for given datasets and integrating these preprocessing capabilities seamlessly into the existing multi-agent system are essential steps.
@@ A.1 Pattern Selection and Weight Optimization for Score Integration
$S_{\text{merged}}=1.2S_{\text{AI}}+2.2S_{\text{KG}}+0.5S_{\text{Search}}$ (7)

### v3

@@ Abstract
Drug repurposing, the identification of new therapeutic applications for existing drugs, offers a promising approach to accelerate drug development. This paper introduces a novel multi-agent framework that enhances the drug repurposing process by leveraging state-of-the-art machine learning techniques and integrating diverse knowledge sources. Our framework consists of three specialized agents: an AI Agent that develops robust drug-target interaction (DTI) models; a Knowledge Graph Agent that extracts DTIs from established databases such as DGIdb, DrugBank, CTD, and STITCH; and a Search Agent that engages with biomedical literature to annotate and verify computational predictions. By synthesizing outputs from these agents, our system harnesses a wide array of data sources to propose viable repurposing candidates. Preliminary results demonstrate the framework’s potential to not only predict drug-disease interactions but also to reduce the time and costs associated with traditional drug discovery methods. Our approach outperforms existing methods in predicting drug repurposing potential while providing interpretable results. This work highlights the scalability of multi-agent systems in biomedical research and their capacity to drive innovation in drug repurposing. The framework presented in this paper offers a promising pathway toward more efficient and cost-effective drug discovery processes, paving the way for accelerated therapeutic developments. Code is available https://anonymous.4open.science/r/DrugRepurposeLLM-E0B5/.
@@ 1 Introduction
Our framework includes three primary agents: the AI agent, the Knowledge Graph (KG) Agent, and the Search Agent. The AI agent employs the DeepPurpose package (Huang et al., 2020) to develop and refine drug-target interaction (DTI) models. The KG Agent utilizes the drug-gene interaction database (DGIdb) (Cannon et al., 2024), DrugBank (Knox et al., 2024), Comparative Toxicogenomics Database (CTD) (Davis et al., 2023), and Search Tool for Interactions of Chemicals (STITCH) (Kuhn et al., 2007) to extract and synthesize information on DTIs. The search Agent engages with biomedical literature, leveraging large language models for automated data labeling and validation, a process resource-intensive and costly.
@@ 2 Related Works
Machine Learning in Drug Repurposing. Machine learning techniques have been applied to drug repurposing, demonstrating significant potential in predicting drug-disease interactions (Issa et al., 2021). Similarly, our AI Agent leverages the DeepPurpose toolkit, which has been validated for its efficiency in drug-target interaction (DTI) predictions (Huang et al., 2020).
@@ 3.1 Overview of DrugAgent
To process natural language inputs and generate responses that are coherent and appropriate, each agent utilizes large language models. The system’s reasoning capabilities are enhanced by incorporating methodologies that allow for step-by-step problem-solving and decision-making.
Following the reasoning process, the system is capable of taking actions such as calculating repurposing scores, performing literature-based analyses, and querying knowledge graphs for novel drug-disease associations. By integrating this information through a weighted average approach, the system simulates a knowledgeable drug repurposing researcher.
@@ 3.2 Agent Roles and Responsibilities
The DrugAgent framework integrates a diverse array of specialized agents, each employing the ReAct (Yao et al., 2022) and LEAST-TO-MOST (Zhou et al., 2022) reasoning methods to plan their actions. Through the use of advanced search capabilities, access to specialist models, and indexing in databases, these agents can execute a wide range of tasks effectively. Below, we delve into the specific roles and responsibilities assigned to each agent within the system.
@@ 3.3 AI Agent
The AI Agent’s predictions are refined through an iterative training and validation process, using cross-validation techniques to ensure accuracy and robustness.
@@ 3.5 Search Agent: Information Extraction from Biomedical Literature
This simplified approach allows for rapid information gathering from publicly available sources. However, it is important to note that this method relies on the quality and relevance of Google search results, and does not analyze full scientific papers or curated databases. As such, it serves as a preliminary screening tool rather than a comprehensive literature review system.
@@ 3.6 Calling External Tools
GPT supports calling external tools (e.g., function, database retrieval) to leverage external knowledge and enhance its capability. Specifically, suppose we have a couple of toolkits, GPT’s API can detect which tool to use, which serves as glue to connect large language models to external tools. Our system integrates a variety of external data sources and predictive AI models to support the agents’ functions.
@@
  • Comparative Toxicogenomics Database (CTD): The CTD (Davis et al., 2023) is a robust, curated database that provides information about chemical–gene/protein interactions, chemical–disease, and gene-disease relationships. It is valuable for understanding how environmental exposures affect human health, integrating data from various species and linking chemicals, genes, diseases, phenotypes, and pathways (Chang et al., 2019; Wu et al., 2022a).
@@
We utilize DeepPurpose (Huang et al., 2020) for the AI Agent. DeepPurpose is a comprehensive and extensible deep learning library for DTI prediction. It integrates multiple state-of-the-art models and datasets, allowing researchers to implemen various deep learning approaches for drug discovery and repurposing. DeepPurpose facilitates the application of AI in drug development by providing a unified framework for different drug and protein encoding methods.
@@ 3.7 Drug Repurposing Agent
The predictions from the AI agent are cross-referenced and enriched with data from the KG Agent and validated against findings from the Search Agent. This integrated workflow allows for a dynamic updating mechanism, wherein feedback from the literature and knowledge graph continually refines the predictions. The entire process is supported by a central data hub that manages data consistency and real-time updates, ensuring that all agents have access to the latest and most accurate data available.
@@
$S_{merged}=\alpha S_{AI}+\beta S_{KG}+\gamma S_{Search},$
where $S_{merged}$ is the merged DTI score, $S_{AI}$ is the AI-based DTI score, $S_{KG}$ is the knowledge graph-based DTI score, $S_{Search}$ is the search-based DTI score, $\alpha$, $\beta$, and $\gamma$ are the weights assigned to the AI, knowledge graph, and search-based scores, respectively.
@@ 3.8 Workflow
Step 6: Delivery of Solution. The final output, which includes the merged DTI score along with the individual scores from each method, is delivered to the user. This comprehensive result provides not only the final prediction but also insights into how different methods contribute to the overall score, enhancing the user’s understanding of the DTI prediction. This structured workflow ensures that our multi-agent DTI prediction system combines multiple analytical approaches, offering a robust and multi-faceted assessment of potential DTIs.
@@ 4.1 Quantitative Results: Performance Comparison of DrugAgent and GPT-4
Table 1: Performance comparison between DrugAgent and GPT-4 for different datasets. We report the average results of 5 independent runs and the corresponding standard deviations (in brackets). For each metric, we highlight the best method in bold. We marked the metrics where DrugAgent is better than GPT-4 (pass the t-test, i.e., p-value$<$0.05) using “*”.
@@
DrugAgent outperformed GPT-4 across all examined metrics. Regarding prediction accuracy, DrugAgent achieved a Mean Squared Error (MSE) of 1.836, lower than GPT-4’s 13.420, indicating superior overall predictive power. The Mean Absolute Percentage Error (MAPE) and Mean Absolute Error (MAE) metrics further confirm this superiority, with DrugAgent achieving values of 0.134 and 1.081, respectively, compared to GPT-4’s 0.320 and 3.350.
@@
These results showcase DrugAgent’s improved predictive capabilities for BindingDB PKD scores compared to GPT-4, which is crucial for accurate binding affinity predictions in drug discovery and molecular interaction studies.
@@
DrugAgent achieves its superior performance with a runtime of 5.000s, which is efficient considering the complexity of its multi-agent architecture. In comparison, GPT-4’s runtime of 0.297s is faster, but at the cost of reduced accuracy.
@@
This comprehensive analysis highlights the synergistic interaction between the AI, KG, and Search Agents in our model architecture for accurate PKD score prediction. The results validate our design choices and provide valuable insights for future improvements and optimizations of DrugAgent. They emphasize the potential for enhancing the integration and capabilities of each component to achieve even more accurate PKD score predictions in the context of the BindingDB database, while considering the balance between performance, computational resources, and cost.
@@ 4.2 Case Study
This study presents three case studies analyzing the drug repurposing potential of Topotecan for different proteins: TOP1, SLFN11, and SLC26A4. These cases represent a spectrum from known strong interactions to potentially novel connections, allowing for a comprehensive evaluation of our multi-agent system’s capabilities. In each case, the Coordinator Agent decomposed the complex problem into three subproblems, each handled by a specialized agent. The AI Agent predicted interaction potential based on structural properties, the Search Agent analyzed evidence from scientific literature, and the KG Agent evaluated connections in existing databases.
@@
Case 1 examined the interaction between Topotecan and TOP1, a known strong interaction. The system calculated a final score of 11.51, confirming the established relationship. The high AI Agent score (7.65) and KG Agent score (1.0) aligned with the known mechanism of Topotecan as a TOP1 inhibitor, while the relatively low Search Agent score (0.27) likely reflected the well-established nature of this interaction not requiring extensive new studies.
@@
Case 2 investigated the less understood but potentially relevant interaction between Topotecan and SLFN11. The system yielded a final score of 10.30, suggesting a noteworthy relationship. The high AI Agent score (7.36) indicated structural compatibility, while the moderate Search Agent (0.33) and KG Agent (0.72) scores reflected some existing evidence and established connections, albeit less comprehensive than in Case 1.
@@
Case 3 explored an unlikely interaction between Topotecan and SLC26A4, resulting in a final score of 9.92. Despite a high AI Agent score (7.61) suggesting structural compatibility, the very low Search Agent score (0.00) indicated minimal literature evidence. The moderate KG Agent score (0.72) suggested some indirect connections, highlighting the system’s ability to detect potential novel interactions while also recognizing the need for careful interpretation.
@@
These case studies demonstrated the multi-agent system’s ability to handle diverse scenarios, from known interactions to potential discoveries. The system integrated various data sources and analytical methods, providing interpretable results with detailed reasoning processes. Importantly, it showed the capacity to recognize its own limitations, particularly when predictions conflicted with existing evidence.
@@
The system offered practical insights for drug repurposing research across different levels of prior knowledge. In Case 1, it confirmed a well-established interaction. In Case 2, it suggested a potentially relevant interaction that warrants further investigation. In Case 3, it identified a possible novel interaction while flagging the low literature evidence, thus highlighting an area requiring careful experimental validation.
@@
This comprehensive approach not only validates known interactions but also suggests potential new avenues for research. By providing nuanced analyses that consider multiple factors, the system demonstrates its value in accelerating and refining the drug repurposing process. Future work could focus on experimental validation of the system’s predictions for cases like Topotecan-SLFN11 where moderate evidence suggests a promising avenue for investigation.
@@ 5 Discussion
Our study presents a novel multi-agent system for drug repurposing that integrates machine learning, knowledge graphs, and literature search. This approach offers more robust predictions by leveraging diverse data sources and analytical methods. The system’s strength lies in its collaborative approach, which combines each agent’s specialized capabilities to evaluate complex DTIs comprehensively. The weighted integration method allows for flexible adjustment of different prediction methods, enhancing overall accuracy. However, limitations exist. The system still relies on human expertise for initial setup, limiting its scalability. It also lacks autonomous knowledge updating capabilities to keep pace with evolving pharmacological research. Furthermore, the current model does not account for individual patient characteristics or drug combination effects.
@@
Automating the preprocessing of complex data types, such as scRNA-seq data, is another important area for improvement. This involves implementing multiple preprocessing functions including Imputation (Inoue et al., 2024a; Inoue, 2024), Quality Control (McCarthy et al., 2017; Lu et al., 2023), and batch effect correction (Li et al., 2020; Fu et al., 2024; Haghverdi et al., 2018). Developing an optimization framework to automatically select and apply the most appropriate preprocessing methods for given datasets and integrating these preprocessing capabilities into the existing multi-agent system are essential steps.
@@ A.1 Pattern Selection and Weight Optimization for Score Integration
$S_{\text{merged}}=1.2S_{\text{AI}}+2.2S_{\text{KG}}+0.5S_{\text{Search}}.$ (7)
While initial optimization suggested a negligible contribution from the Search agent, we assigned it a small but non-trivial weight of 0.5. This decision maintains model flexibility and acknowledges the potential value of diverse information sources in future applications or different datasets, even if not significantly impactful in our current study.